OpenStudy (anonymous):

Use points A= (12,-5) and B= (-4,3) a) find the distance between the points b) find the midpoint

OpenStudy (anonymous):

Do you know the distance formula?

OpenStudy (anonymous):

sorta but not really.

OpenStudy (anonymous):

\[\sqrt{(x2 - x1)^{2} + (y2-y1)^{2}}\]

OpenStudy (anonymous):

That is your formula you pick two points: (x1, y1) and (x2, y2) Plug in that to the formula

OpenStudy (anonymous):

okay! thanks!

OpenStudy (anonymous):

use distance formula in the first part ( as mentioned by @OrthodoxMan ) and in the second part use the formula \[x = (x _{1} + x _{2})/ 2\] and \[y = (y _{1} + y _{2})/2 \] and in this equation x1,x2 ,y1 and y2 are the co-ordinates of the given points :)

OpenStudy (anonymous):

thanks guys!

OpenStudy (anonymous):

Here is my first part...am I correct? \[\sqrt{(12+4)^2 + (-5-3)^2} = \sqrt{(16)^2} = \sqrt{32 + 16} = \sqrt{48} = 6.928?\]

OpenStudy (anonymous):

For part b: I got x=4 and y=-1...so how do I write my answer to the question for part b?

OpenStudy (anonymous):

@Shizen @OrthodoxMan

OpenStudy (anonymous):

for part b, it's just the coordinates (x,y)

OpenStudy (anonymous):

Shizen, you answer part a

OpenStudy (anonymous):

@djibben615 you can't add the squares that ways! cos \[a ^{2} + b ^{2} = (a+b)^{2} - 2ab\]

OpenStudy (anonymous):

in other words Find out the sqaure of 16 and the square of 8 and then add them and find out their square root

OpenStudy (anonymous):

\[\sqrt{256 + 64 } = ?\] find this question mark and thats your answer :)

OpenStudy (anonymous):

And now lemme tell you about the mistake that you've been making you added the 16^2 + 8^2 in a 'wrong' way :)

OpenStudy (anonymous):

so the answer would be 24?

OpenStudy (anonymous):

No! BTW, can you show me the procedure of getting 24..

OpenStudy (anonymous):

look square the 16 which would be 16*16 = 256 and square 8 which would be 8*8 = 64 and now add these 256 + 64 = 320 and now get its square root :) which would be 17 something

OpenStudy (anonymous):

and i repeat it again \[16^{2} + 8^{2} \neq 24^{2}\] and \[16^{2} + 8^{2} = (16+8)^{2} - 2 (16)(8)\]

OpenStudy (anonymous):

ah!!! LIGHTBULB! so its 17 then?

OpenStudy (anonymous):

what about part 2? is the answer: x=4 and y=-1?

OpenStudy (anonymous):

17.88 ! i.e. square root of 320

OpenStudy (anonymous):

yup :) you are correct with the second :)

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Up to helping with another problem? It is attached...

OpenStudy (anonymous):